OpenStudy (anonymous):

A vending machine offers 8 different drinks. One day, 6 employees each purchased a drink from the vending machine. Find the probability that at least 2 employees purchased the same drink. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

each employee has 8 choices of drink, so total # of ways they can choose = 8^6 ways in which all take different drinks = 8P6 = 8*7*6*..*3 so ways in which at least 2 take the same drink = 8^6 - 8P6 probability = (8^6 - 8P6) /8^6 = .9231

OpenStudy (kropot72):

This can be found from the binomial distribution: The probability of choosing a particular drink is 1/8. \[P(0\ out\ of\ 6\ select)=(\frac{7}{8})^{6}=0.4488\] \[P(1\ out\ of\ 6\ select)=6\times \frac{1}{8}\times(\frac{7}{8})^{5}=0.3847\] The probability that at least 2purchase the same drink is 1 - (0.4488 + 0.3847)= ?